### v1

# LoRA is All You Need for Safety Alignment of Reasoning LLMs

Abstract
Reasoning LLMs have demonstrated remarkable breakthroughs in solving complex problems that were previously out of reach. To ensure LLMs do not assist with harmful requests, safety alignment fine-tuning is necessary in the post-training phase. However, safety alignment fine-tuning has recently been shown to significantly degrade reasoning abilities, a phenomenon known as the “Safety Tax”. In this work, we show that using LoRA for SFT on refusal datasets effectively aligns the model for safety without harming its reasoning capabilities. This is because restricting the safety weight updates to a low-rank space minimizes the interference with the reasoning weights. Our extensive experiments across four benchmarks covering math, science, and coding show that this approach produces highly safe LLMs—with safety levels comparable to full-model fine-tuning—without compromising their reasoning abilities. Additionally, we observe that LoRA induces weight updates with smaller overlap with the initial weights compared to full-model fine-tuning. We also explore methods that further reduce such overlap—via regularization or during weight merging—and observe some improvement on certain tasks. We hope this result motivates designing approaches that yield more consistent improvements in the reasoning–safety trade-off.
1 Introduction
Large language models (LLMs) have made remarkable progress across a wide range of tasks. A major recent breakthrough is the emergence of LLMs with advanced reasoning capabilities, enabling them to solve complex problems previously out of reach. However, recent studies have reported significant safety risks associated with reasoning-capable models (Jiang et al., 2025; Zhou et al., 2025; Huang et al., 2025; Li et al., 2025a). Indeed, reasoning fine-tuning—the process through which LLMs acquire these capabilities—often compromises safety, even when starting from a safety-aligned checkpoint (Jiang et al., 2025; Zhou et al., 2025; Zhao et al., 2025; Li et al., 2025a). For example, Jiang et al. (2025) show that models distilled for reasoning from DeepSeek-R1 become substantially less safe than their original base models.
There has been significant effort in the literature to preserve LLMs’ safety alignment during instruction fine-tuning. However, these approaches are not applicable to reasoning fine-tuning. First, reasoning fine-tuning datasets are often highly curated (Muennighoff et al., 2025) and unlikely to contain unsafe content. Thus, data filtering methods such as those proposed by Shen et al. (2024); Choi et al. (2024); Bianchi et al. (2023) are not applicable. In addition, methods that restrict model updates during fine-tuning (Hsu et al., 2024; Mukhoti et al., 2023) are ineffective in the reasoning setting, as acquiring reasoning capabilities typically requires longer training and more substantial weight updates compared to instruction fine-tuning. To the best of our knowledge, the current literature does not offer any method for safety alignment of reasoning models.
Instead, the prevailing strategy is to apply a secondary safety alignment phase after reasoning capabilities have been acquired. This phase—often implemented via supervised fine-tuning (SFT) or reinforcement learning (RL)—has become a standard step in modern LLM development. Although safety alignment fine-tuning can significantly improve model safety, it often comes at a steep cost to reasoning performance—a phenomenon referred to as the “Safety Tax” (Huang et al., 2025). Even incorporating chain-of-thought (CoT) style reasoning into safety fine-tuning datasets (Jiang et al., 2025) cannot succeed in fully preserving reasoning abilities (Huang et al., 2025).
In this work, we investigate the algorithmic factors that contribute to this trade-off. Existing evidence suggests that safety-related behavior in LLMs is often governed by a small number of dominant directions—either in activation space, such as steering vectors (Panickssery et al., 2023) or refusal features (Arditi et al., 2024; Yu et al., 2024), or in weight space. In particular, Jain et al. (2024); Wei et al. (2024) show that safety-critical weights tend to lie in a low-rank subspace. In our analysis, we find that the model undergoes relatively high-rank changes during full-model fine-tuning (see Figure 1), which results in Safety Tax. This highlights a key insight: although achieving safety may require modifying weights only along a low-rank subspace, full-model fine-tuning permits arbitrary updates, potentially introducing many unnecessary changes that interfere with reasoning.
Our extensive experiments reveal the surprising effectiveness of a simple recipe for safety alignment of reasoning models: Applying LoRA during SFT using a straightforward direct refusal data set. Despite its simplicity, this approach achieves safety performance on par with full-model alignment, while preserving reasoning capability close to that of the original reasoning-tuned model. This result holds for both 7B and 14B models and is validated across four benchmarks spanning mathematics, science, and code generation. It represents a rare “one stone, three birds” outcome: strong safety, strong reasoning, and computational efficiency. Moreover, we find that the performance is highly robust to the hyperparameters and configurations of LoRA.
Next, we explore the weight structure imposed by LoRA to understand the differences it introduces. We find that LoRA updates are not only low-rank by design but also exhibit smaller alignment with the weights of the original reasoning model, compared to those from full-model fine-tuning—in most layers. While the reduction in alignment is small, it may suggest that LoRA updates are less disruptive to reasoning-related weights. We further explore whether explicitly reducing such overlap—via regularization or post-hoc weight merging—can improve safety or reasoning capabilities. We observe mixed results, with one post-hoc method achieving a slightly better reasoning–safety trade-off on the AIME and GPQA datasets. More effort is still needed to develop approaches that lead to more consistent improvements across tasks, which we consider a valuable direction for future work.
2 Related Work
To develop LLMs that are both safe and capable, models can be safety aligned before or after fine-tuning.
Fine-tuning a safety-aligned model. Fine-tuning a safety-aligned model often leads to safety degradation. For instruction fine-tuning, safety degradation is shown across various model architectures and optimization strategies, including full-model and LoRA fine-tuning (Qi et al., 2023; Hsiung et al., 2025; Zhan et al., 2023). To mitigate this issue, several methods have been proposed. Shen et al. (2024); Choi et al. (2024) focus on data filtering, aiming to remove unsafe examples from the fine-tuning data. Bianchi et al. (2023) shows that injecting just a few hundred safety examples during instruction fine-tuning can improve safety. Peng et al. (2025) leverages an existing guardrail model to encourage safe response segments while suppressing unsafe ones. Lyu et al. (2024) emphasizes the importance of prompt templates in preserving safety. Algorithmic approaches have also been explored: Hsu et al. (2024) propose projecting LoRA updates into a “safety subspace” derived from differences between aligned and unaligned models, while Mukhoti et al. (2023) introduce regularization techniques to constrain changes in intermediate representations during fine-tuning.
This approach is, however, not applicable to reasoning models. This is because acquiring reasoning capabilities typically requires longer training and more substantial weight updates compared to instruction fine-tuning. This results in losing the initial safety alignment of the model. Indeed, several recent studies have reported significant safety risks associated with reasoning-capable models Jiang et al. (2025); Zhou et al. (2025); Huang et al. (2025); Li et al. (2025a).
Safety alignment after fine-tuning. Aligning a fine-tuned model is typically done using supervised fine-tuning (SFT) and/or reinforcement learning (RL) (Wei et al., 2021; Griffith et al., 2013; Dai et al., 2023; Ouyang et al., 2022; Rafailov et al., 2023; Bai et al., 2022; Guan et al., 2024). However, this approach introduces a key trade-off: safety alignment can substantially impair model performance. Huang et al. (2025) characterizes this as the “Safety Tax,” showing that safety-aligned models often lose much of their reasoning ability. Jiang et al. (2025) attempted to address this issue by constructing a safety fine-tuning dataset with long chain-of-thought (CoT) responses, but the resulting models still showed noticeable drops in reasoning performance (Huang et al., 2025).
In our work, we show that a simple application of LoRA for safety alignment of reasoning models can effectively align reasoning-capable models without compromising their performance.
3 LoRA for Safety Alignment Without Compromising Reasoning
In this section, we investigate whether the “Safety Tax” (Huang et al., 2025) can be mitigated. In particular, we aim to answer the following question: can we align a model for safety without compromising its reasoning capabilities? We focus on the same setup as in (Huang et al., 2025), where SFT is performed on safety datasets that provide harmful requests paired with refusal responses.
Figure 1: We compute the stable rank of the difference between the full-model fine-tuned model’s weights and those of the original DeepSeek-R1-Distill-Qwen-14B for each layer. Here, the colors indicate the module types, and the x-axis shows the layer indices. We observe that the stable rank is quite high—ranging from around 40 to 100 for most layers
Our key observation is that during full-model fine-tuning, which is used in (Huang et al., 2025), the weights undergo relatively high-rank changes. As shown in Figure 1, we observe high stable ranks for the weight updates, i.e., differences between the fine-tuned model’s weights and those of the initial model, in most layers. However, prior evidence suggests that safety behavior in LLMs is typically governed by only a single or a few directions in the activations (Panickssery et al., 2023; Arditi et al., 2024) and weights (Wei et al., 2024; Jain et al., 2024), indicating that a small low-rank modification may be sufficient to induce safe behavior, without altering the entire weight space. Thus, we conjecture that the degradation in reasoning performance is caused by full-model fine-tuning introducing unnecessary changes in many directions, which interfere with critical weights responsible for reasoning.
To address this, we consider Low-Rank Adaptation (LoRA) (Hu et al., 2022), originally proposed as a parameter-efficient fine-tuning method to reduce training cost and memory usage. Rather than updating the full weight matrices, LoRA injects trainable low-rank matrices into existing layers while keeping the original weights frozen. Formally, a weight matrix ${\bm{W}}\in\mathbb{R}^{d\times k}$ is modified as:
${\bm{W}}^{\prime}={\bm{W}}+\Delta{\bm{W}},~{}~{}~{}~{}\text{where}~{}~{}~{}~{}\Delta{\bm{W}}=\frac{\alpha}{r}{\bm{B}}{\bm{A}},$ (1)
where ${\bm{B}}\in\mathbb{R}^{d\times r}$ and ${\bm{A}}\in\mathbb{R}^{r\times k}$ are the trainable low-rank matrices with $r\ll\min(d,k)$, and $\frac{\alpha}{r}$ is the scaling factor, with $\alpha$ being a hyperparameter.
LoRA is particularly well-suited to our needs: it restricts updates to a low-rank subspace, thereby significantly reducing interference with the original weights. We will show in our experiments that this method works excellently, enabling the model to become safe while maintaining strong reasoning performance across benchmarks. As an additional benefit, LoRA is significantly more computationally efficient than full-model fine-tuning.
4 LoRA Bypasses the “Safety Tax”
In this section, we first introduce our safety fine-tuning and evaluation pipeline. Then, we evaluate models’ safety alignment and reasoning performance after full-model and LoRA safety fine-tuning.
4.1 Safety Alignment Fine-Tuning of Reasoning LLMs
We begin with a reasoning-capable language model and perform safety alignment fine-tuning. Specifically, we apply supervised fine-tuning (SFT) on a safety dataset consisting of harmful questions paired with refusal responses, aiming to teach the model to reject harmful requests. We choose SFT over reinforcement learning (RL)-based techniques because it is simpler, less expensive, and does not require additional components such as a reward model. However, we expect our results to apply to RL safety alignment as well.
In our training setup, we compare two approaches: (1) full-model fine-tuning, as in (Huang et al., 2025), where all model parameters are updated using standard gradient-based optimization; and (2) LoRA fine-tuning, as described in Section 3.
4.2 Evaluation of the Fine-Tuned Model
After safety alignment fine-tuning is completed, we evaluate two aspects of the model: (1) safety, which is assessed using a dataset of harmful questions. We sample responses from the model for these questions and use Llama-Guard-3-8B—a model specialized in safety evaluation and shown to be the strongest safety evaluator in (Jiang et al., 2025)—to determine whether the responses are safe. The safety score is defined as the proportion of questions for which the model’s response is judged to be harmful. (2) reasoning ability, evaluated using multiple standard benchmark datasets containing questions on math, science, and coding—widely used to assess models’ reasoning capabilities. We consider the commonly used metric Pass@1 to measure accuracy on these benchmarks. For each question, we sample $n$ responses, compute the fraction of correct responses, and then average this accuracy over all questions. We set $n=8$.
4.3 Datasets and Models
Models. We conduct experiments on two widely used open-weight reasoning-capable models: DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-14B. Safety evaluation is performed using Llama-Guard-3-8B, which was found to be the most accurate evaluator in (Jiang et al., 2025).
Safety fine-tuning dataset. We use the DirectRefusal dataset, adapted from (Rosati et al., 2024) by (Huang et al., 2025), which provides harmful requests paired with refusal-style answers.
Safety evaluation dataset. We adopt the StrongREJECT dataset (Souly et al., 2024), which consists of 310 policy-violating queries designed to test whether the model behaves safely.
Reasoning benchmarks. We evaluate the models performance on (1) American Invitational Mathematics Examination 2024 (AIME), (2) GPQA (Rein et al., 2024) evaluate mathematical and scientific reasoning, respectively, (3) HumanEval (Chen et al., 2021) and (4) MBPP (Austin et al., 2021) are code generation benchmarks. We also consider the augmented versions created by EvalPlus (Liu et al., 2023), denoted as HumanEval+ and MBPP+.
Training Setup. Full-model fine-tuning is performed for 5 epochs, while LoRA fine-tuning is run for 10 epochs. We save and evaluate checkpoints at every epoch. Unless otherwise stated, LoRA is applied only to the MLP layers with rank $r=4$. In the later part of Section 4.4, we investigate the effect of varying $r$ and applying LoRA to different modules (e.g., MLP vs. attention).
Additional experimental details are deferred to Appendix A.
4.4 LoRA is All You Need for Safety Alignment of Reasoning LLMS
(a) DeepSeek-R1-Distill-Qwen-7B
(b) DeepSeek-R1-Distill-Qwen-14B
Figure 3: LoRA bypasses the “Safety Tax”, achieving safety comparable to that of the full-model fine-tuned model and reasoning performance comparable to the original reasoning model. We plot reasoning performance—measured by Pass@1—against safety scores for different models. For the fine-tuned models, we report results for checkpoints at all epochs. Results on the base versions of HumanEval and MBPP are provided in Figure 8 in the Appendix, where the same patterns hold, but with higher accuracy.
Figure 3 compares the safety and reasoning capabilities at different checkpoints (i.e., epochs), during full-model and LoRA safety alignment fine-tuning. We observe that the base model before safety fine-tuning exhibits high accuracy but low safety. On the other hand, the fully fine-tuned models achieve good safety at the cost of reduced accuracy. In contrast, the LoRA fine-tuned models maintain strong performance in both safety and reasoning (as evidenced by the red points in the upper-right corner of the plot).
For the 7B models, the best LoRA checkpoints are on par with, or even slightly outperform, the base model on three reasoning tasks (AIME, HumanEval+, and MBPP+), while showing a slight underperformance on GPQA. In terms of safety, they fall slightly short of the fully fine-tuned models, with approximately a 0.03 drop in safety score. For the 14B models, there is a persistent but small drop in reasoning accuracy compared to the base model, while the safety performance remains comparable to that of the fully fine-tuned model.
(a) Reasoning performance vs. $r$
(b) Safety score v.s. $r$
(c) Applying LoRA to different modules
Figure 4: In (a) and (b), we show reasoning and safety performance at different LoRA ranks $r$ for the 14B model, respectively. Full-model fine-tuning is included as the rightmost point for reference. Reasoning performance decreases as $r$ increases, while safety first decreases and then increases. Overall, very low ranks (e.g., $r=1,4$) are recommended to achieve strong performance in both reasoning and safety. In (c), we compare applying LoRA to different modules. The lightness of the points indicates the number of fine-tuning epochs, with lighter colors corresponding to earlier epochs. Overall, we observe that the performance is quite similar across the two setups.
Effect of $r$. We vary the LoRA rank $r$ in the range $\{1,4,8,64\}$ for fine-tuning the 14B model, and report results at the final checkpoint in Figures 4(a) and 4(b) (with results across all epochs shown in Figure 9). We also include full-model fine-tuning as the rightmost point, since it allows for full-rank updates. From both the final checkpoint and the full training trajectories, we observe that reasoning performance generally declines as $r$ increases, although the drop is very minor between $r=1$ and $r=8$. This suggests that reasoning performance is not very sensitive to the exact choice of $r$, as long as it remains in a reasonable range. For safety scores, we observe a notable drop when $r$ increases from $4$ to $64$, whereas full-model fine-tuning yields a better safety score than $r=64$. We suspect this may be due to optimization difficulty at intermediate high ranks, while very low-rank (extremely under-parameterized) or full-rank (heavily over-parameterized) setups may benefit from easier optimization dynamics, leading to better safety outcomes. Overall, to achieve strong performance in both reasoning and safety, using a very small $r$ is recommended.
Effect of applying LoRA to different modules. Next, we examine the effect of applying LoRA to different modules. Specifically, we compare applying LoRA only to the MLP layers versus applying it to both the MLP and self-attention layers (i.e., the Q, K, V, and O projections). As shown in Figure 4(c), there is no clear difference in performance between the two setups. We note that applying LoRA only to the MLP layers is more computationally efficient.
5 Exploring the Structure of LoRA Weights
Next, we investigate the alignment between LoRA and reasoning weights and study if reducing the alignment further boosts the performance.
(a) $\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$
(b) $\frac{\|{\bm{U}}_{16}{\bm{U}}_{16}^{\top}\Delta{\bm{W}}\|}{\|\Delta{\bm{W}}\|}$
(c) $\frac{\|{\bm{W}}_{I}\Delta{\bm{W}}^{\top}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$
(d) $\frac{\|{\bm{V}}_{16}{\bm{V}}_{16}^{\top}\Delta{\bm{W}}^{\top}\|}{\|\Delta{\bm{W}}\|}$
Figure 5: LoRA updates exhibit smaller overlap with the original weights compared to the full-model fine-tuning updates. Although the reduction in overlap is sometimes small, it can be observed across most layers for all four metrics, which cover both the column (a)(b) and row spaces (c)(d). The 14B models are used here.
5.1 LoRA Updates Have Less Alignment with Initial Weights
Intuitively, if the initial (reasoning) weights ${\bm{W}}_{I}$ and the LoRA update $\Delta{\bm{W}}$ have only a small alignment, this suggests minimal interference between the safety-oriented update and the weights critical for reasoning. LoRA already constrains $\Delta{\bm{W}}$ to be low-rank, meaning it spans only a small subspace of the full weight space. We further examine the orientation of this subspace: how do the directions learned by LoRA compare to those spanned by ${\bm{W}}_{I}$? To quantify this, we compute the following four metrics: (1) $\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$, (2) $\frac{\|{\bm{U}}_{16}{\bm{U}}_{16}^{\top}\Delta{\bm{W}}\|}{\|\Delta{\bm{W}}\|}$, (3) $\frac{\|{\bm{W}}_{I}\Delta{\bm{W}}^{\top}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|}$, and (4) $\frac{\|{\bm{V}}_{16}{\bm{V}}_{16}^{\top}\Delta{\bm{W}}^{\top}\|}{\|\Delta{\bm{W}}\|}$. Here, ${\bm{U}}_{16}$ and ${\bm{V}}_{16}$ are matrices containing the top 16 left and right singular vectors of ${\bm{W}}_{I}$, respectively, obtained via truncated SVD. Intuitively, (1) and (2) capture the overlap between ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ in the column space: (1) is a matrix-level analogue of cosine similarity, while (2) measures the normalized projection of $\Delta{\bm{W}}$ onto the top dominant directions of ${\bm{W}}_{I}$. Similarly, (3) and (4) capture overlap in the row space. The column and row spaces correspond to the directions the matrices “write to” and “read from”, respectively. A smaller value in any of these metrics indicates greater orthogonality between ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ in the corresponding space.
We compare the full-rank fine-tuned model with the LoRA fine-tuned model in which both the attention and MLP modules are updated (i.e., corresponding to “QKVO & MLP” in Figure 4(c)), making the two settings more comparable since updates occur in all major modules. We compute the alignment metrics for different module types across layers, average them over layers, and report the results for the 14B models in Figure 5. We observe that LoRA achieves smaller overlap in most modules across the metrics, with a few exceptions. This suggests that, for most weights, ${\bm{W}}_{I}$ and $\Delta{\bm{W}}$ are more orthogonal—both in the column and row spaces—for LoRA than for full-model fine-tuning. In other words, under LoRA fine-tuning, the safety-oriented updates read from and write into subspaces that are more separate from those used by the original reasoning-related weights—more so than in the full-model fine-tuned version. Although the reduction in alignment values is sometimes small, it may still indicate that LoRA updates interfere less with the reasoning-related components of the model, potentially explaining the better preservation of reasoning performance. A more in-depth investigation is needed to fully understand the underlying mechanisms and to develop more precise metrics for capturing this effect—an important direction for future work.
(a) 7B models on AIME and GPQA.
(b) 7B models on HumanEval+ and MBPP+.
(c) 14B models on AIME and GPQA.
(d) 14B models on HumanEval+ and MBPP+.
Figure 6: For each method, we visualize only the points on the Pareto frontier of reasoning performance versus safety score. (a)(c) On AIME and GPQA, OrthoMerge-both appears most promising, as its points are more concentrated in the upper-right corner and are sometimes strictly better than those of vanilla LoRA—especially for the 14B models. (b)(d) However, when we compare it to vanilla LoRA on the coding benchmarks, it slightly underperforms in terms of accuracy. Overall, OrthoMerge-both can occasionally outperform vanilla LoRA in reasoning performance while achieving slightly higher safety scores.
5.2 Exploring Methods That Further Reduce Alignment
Given the observations in Section 5.1, we wondered whether further reducing the overlap between $\Delta{\bm{W}}$ and ${\bm{W}}_{I}$ could lead to even better reasoning performance without compromising safety. This question is particularly worth exploring because, while the results achieved by LoRA in Section 4.4 are strong, they are not perfect—a small performance gap remains compared to the original reasoning model, especially for the 7B model on GPQA and the 14B model on AIME, HumanEval+, and MBPP+.
To test this, we consider two approaches for enforcing stronger orthogonality: regularization during LoRA fine-tuning and enforcing orthogonality during merging LoRA weights, as we discuss next.
Regularization during LoRA training. We add a penalty term to the loss that discourages overlap between ${\bm{W}}_{I}$  [^1] To avoid out-of-memory issues during training caused by the large dimensionality of model weights, in implementation we use a low-rank approximation of ${\bm{W}}_{I}$ instead of the full matrix. and $\Delta{\bm{W}}$, specifically:
  • Regularization-col: $\beta(\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|)})^{2}$ encourages orthogonality in the column space.
  • Regularization-both: $\beta(\frac{\|{\bm{W}}_{I}^{\top}\Delta{\bm{W}}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|)})^{2}+\beta(\frac{\|\Delta{\bm{W}}^{\top}{\bm{W}}_{I}\|}{\|{\bm{W}}_{I}\|\|\Delta{\bm{W}}\|})^{2}$ encourages orthogonality in both column and row spaces.
We tried different values of $\beta$ but found no significant change in the results, so we fix $\beta=1$.
Enforcing Orthogonality During Merging LoRA Weights. Starting with updates obtained from standard LoRA, we modify how they are merged with ${\bm{W}}_{I}$, which we call OrthoMerge. Before applying Equation 1, we preprocess $\Delta{\bm{W}}$ as follows:
  • OrthoMerge-col: $\Delta{\bm{W}}\leftarrow({\bm{I}}-{\bm{U}}_{k}{\bm{U}}_{k}^{\top})\Delta{\bm{W}}$ enforces column-space orthogonality based on the rank-$k$ SVD of ${\bm{W}}_{I}$.
  • OrthoMerge-both: $\Delta{\bm{W}}\leftarrow({\bm{I}}-{\bm{U}}_{k}{\bm{U}}_{k}^{\top})\Delta{\bm{W}}({\bm{I}}-{\bm{V}}_{k}{\bm{V}}_{k}^{\top})$ enforces orthogonality in both column and row spaces. We found that directly applying OrthoMerge-both leads to a drop in safety scores. To mitigate this, we further scale up the orthogonal complement with $\Delta{\bm{W}}\leftarrow\lambda({\bm{I}}-{\bm{U}}_{k}{\bm{U}}_{k}^{\top})\Delta{\bm{W}}({\bm{I}}-{\bm{V}}_{k}{\bm{V}}_{k}^{\top})$ to compensates for the loss in safety. We experiment with different values of $\lambda$ in the range $\{1,1.15,1.75,1.2,1.25\}$. We set $k=64$.
For both approaches, we omit the row-space-only variant, as it showed no significant improvement in our experiments.
In Figures 6(a) and 6(c), we show the reasoning performance on AIME and GPQA against safety. Since we have two evaluation dimensions—safety and reasoning performance—for each method that yields multiple checkpoints (e.g., different epochs for vanilla and regularization fine-tuning, or different hyperparameters for OrthoMerge), we only visualize the Pareto-optimal points to compare the best-performing configurations. We observe that methods addressing both the row and column spaces (“both”) tend to yield better results than those that only operate on the column space (“col”). Among them, OrthoMerge-both appears most promising, with more points concentrated in the upper-right corner—often strictly dominating the points corresponding to vanilla LoRA. Therefore, we additionally evaluate OrthoMerge-both on the coding benchmarks and compare it against vanilla in Figures 6(b) and 6(d). However, in this case, it slightly underperforms. Overall, across tasks, our conclusion is that OrthoMerge-both tends to achieve slightly better safety scores and occasionally better reasoning performance than vanilla LoRA, though the improvements are not consistent. This suggests potential for further improving the safety–reasoning trade-off by encouraging orthogonality, and further investigation is needed to identify approaches that yield consistent gains—making this a valuable direction for future research.
6 Conclusion and Discussion
In this work, we identified a simple solution to the previously observed tension between reasoning and safety in reasoning LLMs. Through extensive experiments, we found that fine-tuning a reasoning model with LoRA on a safety dataset preserves its reasoning abilities, whereas full-model fine-tuning significantly degrades them. Additionally, we observe that LoRA induces weight updates with smaller overlap with the initial weights, compared to full-model fine-tuning. This may help explain the better reasoning performance of LoRA safety fine-tuning. We also explore further reducing this overlap via regularization or during weight merging and observe slight improvements on certain tasks. We hope this can inspire future work to design methods that lead to more consistent improvements. Another promising research direction is to understand and characterize how LoRA naturally produces updates that are less disruptive to the model’s reasoning function. This is particularly intriguing because LoRA only constrains the update rank, without explicitly restricting the direction.
Appendix A Experimental Details
A.1 Details about DirectRefusal
The original DirectRefusal dataset can be downloaded from TianshengHuang/DirectRefusal on Hugging Face. For each harmful request, the corresponding response includes a short thought—“I should not answer this question!”—followed by a refusal answer. We adapted the dataset to ensure it follows the DeepSeek-R1 chat template.
A.2 Details about StrongReject
For each question, we sample the answer using greedy decoding and pass the answer to Llama-Guard-3-8B to judge whether it is safe. Safety is determined by checking whether the substring “safe” or “unsafe” is contained in the evaluator’s output.
A.3 Details about AIME and GPQA
We adapted the codebase from Li et al. (2025b), which builds on top of lm-evaluation-harness Gao et al. (2024), to evaluate models on GPQA and AIME. For AIME, answers are evaluated using Qwen2.5-32B-Instruct as a judge. For GPQA, since it is a multiple-choice benchmark, answers are evaluated via regular expression matching. To improve evaluation accuracy, we made some small adjustments. One change is modifying the prompt to better ensure the model outputs answers in the desired format:
[…]
Question: {{Question}}
Choices:
(A) {{choice1}}
(B) {{choice2}}
(C) {{choice3}}
(D) {{choice4}}
Another adjustment is in the answer extraction logic—we enhanced the original implementation to handle a wider range of answer formats that models produce.
During generation, we set the temperature to 0.6, top_p to 0.95, and the maximum number of generated tokens to 32,768.
A.4 Details about HumanEval and MBPP
We adapted the codebase of EvalPlus Liu et al. (2023). The original implementation included a response prefix, designed for earlier models that did not explicitly support an intermediate reasoning process. This prefix—such as “Below is a Python script with a self-contained function that efficiently solves the problem and passes corresponding tests:”—was prepended to model outputs during generation. However, this practice introduces unfair bias by encouraging all models to directly generate code. It inadvertently benefits the full-model fine-tuned baseline—which would otherwise often refuse to answer—by effectively forcing it to produce code. Conversely, it disadvantages reasoning-aligned models by disrupting the expected format that includes intermediate “thought”, causing them to skip the thinking process entirely. This can result in skewed conclusions. We suspect this explains the abnormally low accuracy of the base reasoning model and the high accuracy of the full-model fine-tuned variant reported in Jiang et al. (2025). To address this issue, we remove the response prefix. We also slightly reword the prompt to better align with reasoning models.
During generation, we set the temperature to 0.6 and the maximum number of generated tokens to 32,768.
A.5 Training Details
For 7B models, full-model fine-tuning is conducted using 4 GPUs with a batch size of 2 per device for 5 epochs. LoRA fine-tuning uses 2 GPUs with a batch size of 2 per device for 10 epochs. We set the LoRA hyperparameters as $\alpha=16$ and lora_dropout = 0.05.
For 14B models, full-model fine-tuning is performed using 8 GPUs with a batch size of 1 per device for 5 epochs. LoRA fine-tuning uses 4 GPUs with a batch size of 2 per device for 10 epochs. We set the LoRA hyperparameters as $\alpha=16$ and lora_dropout = 0.05.
Across all experiments, we use a learning rate of 5e-5 and a weight decay of 1e-4.
(a) DeepSeek-R1-Distill-Qwen-7B
(b) DeepSeek-R1-Distill-Qwen-14B
Figure 8: Results on the base versions of HumanEval and MBPP show the same pattern as in the plus versions shown in Figure 3, but with higher accuracy.
Figure 9: Comparing the results for different values of $r$, the lightness of the colors represents the number of epochs, with lighter shades indicating earlier epochs. Overall, we observe that performance is very similar for $r\in\{1,4,8\}$, while $r=64$ leads to a noticeable drop in reasoning performance on GPQA, as well as worse safety. Overall, we conclude that performance remains strong as long as $r$ is not set excessively large.